When 15.0 mL of a 3.74×10-4M ammonium sulfide solution is combined with 12.0 mL of a 8.12×10-4M silver acetate solution does a precipitate form?  (yes or no) _____

For these conditions the Reaction Quotient, Q, is equal to _____
Answer #1

(NH4)2S Ag Chacoo silver acelete Ammonium sulfide - 1som 3.74X16 M 12.0m2 8.12 x 10 lm Total volume of sol = 15-0 +12-0 mL =(MK4)25 - Q NH4 +52 2:08x1014 [s^3= 2: 08 3 low CH3 Coong CH₂ coo- + Ag+ Cagt] = 3.61X10-4 2 Ag++ s2- ² Agas Ksp of Ag, S = 6